Find two positive numbers whose difference is 6 and the sum of whose squares is 818.
Give the answers in ascending order:
x =

Answer by mananth(16946)

if one number is x
the other will be x+6
x^2 + (x+6)^2 = 818
x^2+x^2+12x+36=818
2x^2+12x-782=0
x^2 +6x - 391=0
x^2+6x+9 = 391+9
(x+3)^2=400
x+3=20
x=17
17, 23